OpenStudy (anonymous):

(sqrt2x+13) - 15 = x

OpenStudy (mertsj):

Add 15 to both sides and then square both sides.

OpenStudy (anonymous):

I'm still kind of confused

OpenStudy (mertsj):

Did you add 15 to both sides?

OpenStudy (anonymous):

Yeah, I did.

OpenStudy (mertsj):

What did you get?

OpenStudy (anonymous):

\[\sqrt{2x+13} = x + 15\]

OpenStudy (mertsj):

What do you get if you square the left side?

OpenStudy (mertsj):

\[(\sqrt{2x=13})^2= ?\]

OpenStudy (mertsj):

Sorry. I meant + 13 not =

OpenStudy (theviper):

\[\large{\sqrt{2x+13}-15=x}\]\[\large{\sqrt{2x+13}=x+15}\]Now squaring both sides we get\[\large{2x+13=x^2+225+30x}\]

OpenStudy (mertsj):

You must not square each side. So now it will look like this: \[(\sqrt{2x+13})^2=(x+15)^2\]

OpenStudy (mertsj):

ok. Too many cooks. Good bye

OpenStudy (theviper):

bye :)

OpenStudy (anonymous):

Thanks for leaving I guess... @Mertsj @TheViper are you leaving too or saying goodbye to the other guy?

OpenStudy (theviper):

no

OpenStudy (anonymous):

Okay haha

OpenStudy (theviper):

\[\large{2x+13=x^2+225+30x}\]\[\large{\implies x^2+30x-2x+225-13=0}\]\[\large{\implies x^2+28x+212=0}\]Now solve it by factoring :)

OpenStudy (anonymous):

You two were both telling me different things. How do I know that yours it right?

OpenStudy (theviper):

We are telling same but concentrate on mine because I m here to help you :)

OpenStudy (mertsj):

It won't factor. You'll have to use the quadratic formula.

OpenStudy (anonymous):

I tried factoring and it didn't work. So I plug x^2 + 28x + 212 into the quadratic formula?

OpenStudy (mertsj):

yes

OpenStudy (theviper):

yup :)

OpenStudy (theviper):

I think it should be \(-14+2i\) \(\&\) \(-14-2i\)

OpenStudy (theviper):

Check ur calculation again :)